# Evaluating ion dynamics through Coulomb and Yukawa interaction potentials in one component strongly coupled plasmas

### v1

@@ Abstract
Atmospheric pressure helium plasmas are investigated employing molecular dynamics simulations at room temperature (300 K) for various ionization fractions ($\chi_{i}=10^{-1}-10^{-5}$) in the strongly coupled regime (ion coupling parameter, $\Gamma_{i}\sim 1-10$). The role of electron screening in ion dynamics and energetics is examined through ion and gas temperatures, mean squared displacement of ions, ion coupling parameter, and radial distribution function of the system by implementing both Coulomb and Yukawa potentials for ionic interactions. It is found that electron screening in the Yukawa potential limits the disorder-induced heating (DIH) mechanism and the coupling strength of the ions significantly for strongly ionized plasmas (ionization fraction, $\chi_{i}$ $\geq 10^{-3}$). Whereas, the ions show a prominent sub-diffusive behavior for the Coulomb potential that is associated with the DIH mechanism. DIH is explained using a simplified model based on energy conservation. Moreover, the screening effect affects the separation distance and arrangement of the ion-neutral pairs for all the values of $\chi_{i}$.
@@ I Introduction
Plasmas at atmospheric pressure have been studied extensively in a wide range of experiments and found their applications in surface processing and biomedical fields [1, 2, 3, 4]. The criticality of their research lies in the fact that it has led to the emergence of a new field known as plasma medicine [5, 6]. These plasmas are often generated in laboratories at room temperatures employing RF sources [7], microwaves [8], DC sources [9], nano-second pulsed discharges [10], and dielectric barrier discharges [11]. Depending on the mode of generation, the value of ionization fraction ($\chi_{i}=n_{i}/(n_{i}+n_{n})$, where $n_{i}$ and $n_{n}$ are the ion and gas number densities, respectively) in the atmospheric pressure plasmas, ranges from $10^{-1}$ to $10^{-8}$, and divides the plasma into strong, moderate and weakly ionized regimes. Based on their ionization fraction, the plasmas show significant heating and strong coupling among the ions after their generation [12, 9, 13], where the ion coupling parameter ($\Gamma_{i}$), which is the ratio of mean ion potential energy to the mean ion kinetic energy becomes $\geq 1$.
In strongly ionized plasmas, the ion heating is assumed to be caused by the disorder-induced heating (DIH) mechanism [12]. DIH arises as the ions tend to move to a lower potential energy state from a disordered state and release kinetic energy in the process. However, for weakly ionized plasmas, the heating is often attributed to the discharge geometry and type of plasma source [9, 13].
@@
For the moderate and weakly ionized plasmas, most of the simulation and theoretical efforts are focused on understanding the plasma chemistry, electron temperature ($T_{e}$), and electron densities ($n_{e}$) for a given plasma geometry and generation scheme [9, 13, 14]. Furthermore, the contributions of intra and inter-species interactions and internal structures of the bulk plasmas in plasma heating and ion coupling strength are not very well understood, even for the strongly ionized plasmas. In this context, we employ molecular dynamics (MD) simulations to investigate the ion and neutral gas dynamics and their implications in ion energies, ion coupling strength, ion and neutral temperatures, and the internal structures in an atmospheric pressure helium plasma for $\chi_{i}$ (= $10^{-1}$ to $10^{-5}$), at room temperature (300 K), thus covering the strong, moderate and weakly ionized plasma regimes. MD simulations efficiently track each particle’s trajectories and easily convey molecular-level information [15, 16, 17, 18]. However, while employing the MD simulations, it is imperative to designate the intra and inter-species interactions with appropriate potentials. Therefore, in this work, we discuss the performance of both Coulomb and Yukawa potentials as the probable ion-ion interaction potential for atmospheric pressure plasmas, which may help us to decide the appropriate potential, depending upon $\chi_{i}$ and other plasma parameters. Section II of this article discusses the inter-particle potentials in-depth, along with the simulation details, followed by the important results outlined in section III and a discussion in section IV. The article is concluded in section V.
@@ III.2 Ion coupling parameter
The mean kinetic energy ($KE$) for both potentials is $k_{B}T_{i}$. $T_{i}$ depends on the chosen potential at the equilibrium condition. Subscripts “C” and “Y” are used to refer to Coulomb and Yukawa potentials, respectively. Figure 4 shows the variation of $PE$ and $KE$ with $\chi_{i}$ for both potentials. It is evident that $PE$ for Yukawa potential is lower than that for Coulomb potential, which provides an energetically more favorable condition for the plasma system. Meanwhile, $KE$ remains almost constant for both potentials except for the $\chi_{i}=10^{-1}$ case.
@@ IV Discussion
To understand the potential energy contribution to the DIH and oscillations in the $T_{i}$ curve, a simple model is developed. According to the DIH hypothesis, the increase in ion kinetic energy is attributed to the decrease in their potential energy. Mathematically, for an individual ion, it can be written as
@@ V Conclusion
It is found that the DIH, which is prominent for $\chi_{i}\geq 10^{-3}$, is more significant for the Coulomb potential. The electron screening in the Yukawa potential efficiently limits the ion and gas heating in the system. During the non-equilibrium phase ($T_{i}\neq T_{g}$) caused by the DIH, the $T_{i}$ curve for the Coulomb potential shows an oscillatory behavior for $\chi_{i}\geq 10^{-3}$. The oscillations are very much damped for the Yukawa potential. For weakly ionized plasmas ($\chi_{i}\leq 10^{-4}$), such oscillations are absent for both potentials. During this non-equilibrium phase, the ions show a sub-diffusive behavior for $\chi_{i}\geq 10^{-3}$, which is more prominent for the Coulomb potential. The sub-diffusive behavior is exclusively associated with the DIH mechanism.
@@
To conclude, it is more scientific to include the screening effect in the simulations for the atmospheric pressure plasma systems, as this effect dictates the plasma dynamics, energetics, and internal structure more realistically. Although the Coulomb potential is easier to implement with long-range effects and is computationally less expensive, the Yukawa potential may be important for ascertaining more physical results for the given experimental conditions. However, the Yukawa potential needs to be investigated further to ascertain how it guides the ion-neutral structural arrangements at room-temperature plasmas and to extend the current finding to high and low gas temperatures at atmospheric pressure, which may be taken up in the future works.

### v2

@@ Abstract
Atmospheric pressure helium plasmas are investigated through molecular dynamics simulations at room temperature (300 K) for various ionization fractions ($\chi_{i}=10^{-1}-10^{-5}$) in the strongly coupled regime (ion coupling parameter, $\Gamma_{i}\sim 1-10$) employing Coulomb and Yukawa interaction potentials. The role of electron screening in ion dynamics and energetics is examined through ion and gas temperatures, mean squared displacement of ions, ion coupling parameter, and radial distribution function of the system. It is found that electron screening in the Yukawa potential significantly limits the disorder-induced heating (DIH) mechanism for strongly ionized plasmas ($\chi_{i}$ $\geq 10^{-3}$). Whereas, ions show a prominent sub-diffusive behavior associated with the DIH during the non-equilibrium phase for the Coulomb potential. The DIH mechanism is explained using a model based upon the conservation of energy. However, for weakly ionized plasmas ($\chi_{i}\leq 10^{-4}$), the maximum ion temperatures are almost similar for both potentials. Furthermore, electron screening affects the separation distance and arrangement of the ion-neutral pairs for all the values of $\chi_{i}$. In general, Yukawa potential results in a lower mean potential energy of the interacting particles, which is energetically favourable for the stability of the system.
@@ I Introduction
Plasmas at atmospheric pressure have been studied extensively in a wide range of experiments and found their applications in surface processing and biomedical fields [1, 2, 3, 4]. The criticality of their research lies in the fact that it has led to the emergence of a new field known as plasma medicine [5, 6]. These plasmas are often generated in laboratories at room temperatures employing RF sources [7], microwaves [8], DC sources [9], nano-second pulsed discharges [10], and dielectric barrier discharges [11]. Depending on the mode of generation, the value of ionization fraction ($\chi_{i}=n_{i}/(n_{i}+n_{n})$, where $n_{i}$ and $n_{n}$ are the ion and gas number densities, respectively) in the atmospheric pressure plasmas, ranges from $10^{-1}$ to $10^{-8}$, and divides the plasma into strong, moderate and weakly ionized regimes. Based on their ionization fraction, the plasmas show significant heating and strong coupling among the ions after their generation [12, 9, 13], where the ion coupling parameter ($\Gamma_{i}$), which is the ratio of mean ion potential energy to the mean ion kinetic energy becomes $\geq 1$. In strongly ionized plasmas, the ion heating is assumed to be caused by the disorder-induced heating (DIH) mechanism [12]. DIH arises as the ions tend to move to a lower potential energy state from a disordered state and release kinetic energy in the process. However, for weakly ionized plasmas, the heating is often attributed to the discharge geometry and type of plasma source [9, 13]. For the moderate and weakly ionized plasmas, most of the simulation and theoretical efforts are focused on understanding the plasma chemistry, electron temperature ($T_{e}$), and electron densities ($n_{e}$) for a given plasma geometry and generation scheme [9, 13, 14]. Furthermore, the contributions of intra and inter-species interactions and internal structures of the bulk plasmas in plasma heating and ion coupling strength are not very well understood, even for the strongly ionized plasmas. In this context, we employ molecular dynamics (MD) simulations to investigate the ion and neutral gas dynamics and their implications in ion energies, ion coupling strength, ion and neutral temperatures, and the internal structures in an atmospheric pressure helium plasma for $\chi_{i}$ (= $10^{-1}$ to $10^{-5}$), at room temperature (300 K), thus covering the strong, moderate and weakly ionized plasma regimes. MD simulations efficiently track each particle’s trajectories and easily convey molecular-level information [15, 16, 17, 18]. However, while employing the MD simulations, it is imperative to designate the intra and inter-species interactions with appropriate potentials. Therefore, in this work, we discuss the performance of both Coulomb and Yukawa potentials as the probable ion-ion interaction potential for atmospheric pressure plasmas, which may help us to decide the appropriate potential, depending upon $\chi_{i}$ and other plasma parameters. Section II of this article discusses the inter-particle potentials in-depth, along with the simulation details, followed by the important results outlined in section III and a discussion in section IV. The article is concluded in section V.
@@ III.2 Ion coupling parameter
The mean kinetic energy ($KE$) for both potentials is $k_{B}T_{i}$. $T_{i}$ depends on the chosen potential at the equilibrium condition. Subscripts “C” and “Y” are used to refer to Coulomb and Yukawa potentials, respectively. Figure 4 shows the variation of mean $PE$ and $KE$ with $\chi_{i}$ for both potentials. It is evident that the mean $PE$ for Yukawa potential is lower than that for Coulomb potential, which provides an energetically more favorable condition for the plasma system. Meanwhile, the mean $KE$ remains almost constant for both potentials except for the $\chi_{i}=10^{-1}$ case.
@@ IV Discussion
To understand the potential energy contribution to the DIH and oscillations in the $T_{i}$ curve, a model is developed. According to the DIH hypothesis, the increase in ion kinetic energy is attributed to the decrease in their potential energy. Mathematically, for an individual ion, it can be written as
@@ V Conclusion
It is found that the DIH, which is prominent for $\chi_{i}\geq 10^{-3}$, is more significant for the Coulomb potential. The electron screening in the Yukawa potential efficiently limits the ion and gas heating in the system. The maximum ion temperatures for the weakly ionized plasmas ($\chi_{i}\leq 10^{-4}$) are almost similar for both potentials. During the non-equilibrium phase ($T_{i}\neq T_{g}$) caused by the DIH, the $T_{i}$ curve for the Coulomb potential shows an oscillatory behavior for $\chi_{i}\geq 10^{-3}$. The oscillations are very much damped for the Yukawa potential. For weakly ionized plasmas ($\chi_{i}\leq 10^{-4}$), such oscillations are absent for both potentials. During this non-equilibrium phase, the ions show a sub-diffusive behavior for $\chi_{i}\geq 10^{-3}$, which is more prominent for the Coulomb potential. The sub-diffusive behavior is exclusively associated with the DIH mechanism.
@@
To conclude, it is more scientific to include the screening effect in the simulations for the atmospheric pressure plasma systems, as this effect dictates the plasma dynamics, energetics, and internal structure more realistically. The mean potential energy in a Yukawa system is lower, which is energetically favourable for the stability of the system. Although the Coulomb potential is easier to implement with long-range effects and is computationally less expensive, the Yukawa potential may be important for ascertaining more physical results for the given experimental conditions. However, the Yukawa potential needs to be investigated further to ascertain how it guides the ion-neutral structural arrangements at room-temperature plasmas and to extend the current finding to high and low gas temperatures at atmospheric pressure, which may be taken up in the future works.